OpenStudy (anonymous):

The regular price of an item is $90. The item is on sale at a discount rate of 40%. What is the sale price of the item?

OpenStudy (anonymous):

36

OpenStudy (anonymous):

90-36=54

OpenStudy (anonymous):

90*0.40=36 90-36=54